A toy is assembled from 3 parts.If the probability that each part is defected is 0.1.Calculate the probability that the toy is defective.
Answer by stanbon(75887)

--------------------A toy is assembled from 3 parts.If the probability that each part is defected is 0.1.Calculate the probability that the toy is defective.
-------------
Probability a part is NOT defective = 0.9
Probability all three parts are NOT defective = 0.9^3 = 0.729
-----------
Prob (the toy is defective)+ Prob (the toy is not defective) = 1
So Prob (the toy is defective) = 1 - 0.729 = 0.271
=================
Cheers,
Stan H.